# UniGame: Turning a Unified Multimodal Model Into Its Own Adversary

### v1

@@ 1 Introduction
Unified Multimodal Models (UMMs) have recently demonstrated impressive capability in both visual understanding and image generation with a unified architecture [deng2025emerging, wang2024emu3, qu2025tokenflow, xie2024show, xiao2025omnigen, wu2025harmonizing]. By jointly leveraging a language model backbone and a visual tokenizer–decoder stack [li2025dual, chen2024janus], these models promise a unified interface for cross-modal reasoning, grounded perception, and controllable generation. Specifically, the large-scale pre-training establishes general multimodal capabilities, and the post-training stage (supervised fine-tuning, SFT, Figure 2(a)) can further improve their performance on downstream tasks with enhanced reliability.
@@
Despite their great performance, UMMs exhibit structural inconsistency between their understanding and generation pathways [yan2025can, wang2024reconstruction]. This inconsistency stems from the inherently conflicting nature of the two objectives, leading to the mismatch in various aspects such as semantics (i.e., the model can answer a question correctly yet fail to generate a corresponding image, or vice versa [Gan2020, rajabi2025token]), capability (e.g., generation is harder to improve than understanding, or vice versa [radford2021learning]), and feature compactness (e.g., understanding requires more compact feature space while generation prefers oppositely). Inconsistency widely exists in real-world applications, where models frequently encounter unexpected inputs far from the training manifold, compositional combinations unseen during training, counterfactual queries, or modality conflicts such as distribution shift [Gan2020, li2024naturalbench, oh2025understanding, oh2025vittle] and adversarial attack [li2021adversarial]. If not sufficiently studied, it would greatly undermine multimodal information fusion, model robustness, and further performance improvement.
@@
It remains challenging and unexplored to improve the consistency of UMMs, primarily due to the unclear learning objectives: only the consequences, but not the causes, are known. Therefore, recent post-training approaches tend to fill this gap using surrogate objectives. Reconstruction-based approaches (Figure 2(b)) regenerate the original images through the semantic embedding space derived from visual perception [wang2024reconstruction, yan2025can]. They optimize a unified reconstruction objective, which trains models within a closed auto-encoding loop. Reward-based methods (Figure 2(c)) typically optimize an ensembled reward function [jiang2025t2i], combining task-specific or rule-based metrics to refine the output relying on external expert models. However, both are optimized using handcrafted objectives (reconstruction or reward), which only polish model behavior on a fixed training distribution and place no explicit constraints on the two coupling branches. As a result, they reproduce behaviors within a fixed manifold rather than expanding the shared generative space, leaving the inconsistency largely unresolved.
@@ 2 Related Work
Post-training of UMMs. Aiming to resolve the inconsistency issue, existing post-training methods can be categorized into these types: (i) reconstruction and semantic-alignment losses to encourage fidelity [wang2024reconstruction, yan2025can]; (ii) RL/reward-based optimization to directly optimize downstream metrics [Ramesh2024, jiang2025t2i]. Each family improves either fidelity or robustness [Li2023, chen2024janus, wang2024reconstruction, Gan2020, rajabi2025token, Ramesh2024, qu2025tokenflow], For instance, RecA [wang2024reconstruction] leverages reconstruction alignment by conditioning generation on understanding embeddings and using reconstruction losses to bring representations closer, while T2I-R1 [jiang2025t2i] enhances image generation through collaborative semantic- and token-level chain-of-thought combined with reinforcement learning. In the vision–language domain, AT has shown potential: VILLA introduces large-scale embedding-space perturbations across image and text modalities, improving robustness and generalization [Gan2020, rajabi2025token]. Nevertheless, most methods either improve alignment or generation separately, and adversarial mechanisms are rarely incorporated into full UMMs. They commonly fail in one crucial respect: they do not exploit generation as an active adversarial process to strengthen the understanding branch. Specifically, adversarial or unguided embedding perturbations often produce off-manifold samples—unrealistic or semantically invalid artifacts, while reconstruction objectives do not intentionally surface decision-critical failure modes. Although RL-based schemes are effective, they are computationally costly and do not guarantee that these discovered examples remain decodable and semantically plausible [Ramesh2024, jiang2025t2i]. This leaves a key gap in truly improving the consistency of the understanding and generation pathways.
@@ 3.2 Motivation
Existing efforts [wang2024reconstruction, jiang2025t2i] optimize for individual goals on a fixed set of data distributions, and place no explicit constraints on the two coupling training objectives. They encourage cooperative training, reproducing existing samples rather than expanding coverage of the shared manifold, where boundary behavior is most fragile, thus aggravating inconsistency. We argue that improving consistency requires expanding this shared manifold, especially around decision boundaries, instead of merely polishing the model within its comfort regions.
@@ 3.4 The Self-Play Training Process
$\begin{split}\mathcal{L}_{U}(\theta_{U})=\;&\mathbb{E}_{\mathrm{clean}}\big[\mathrm{CE}(p_{U}(\hat{a}\mid\mathbf{z},q;\theta_{U}),a)\big]\\ &+\beta\,\mathbb{E}_{\mathcal{B}}\big[\mathrm{CE}(p_{U}(\hat{a}\mid\mathbf{z},q;\theta_{U}),a)\big],\end{split}$ (6)
@@
[…]
@@ 3.5 Discussion
Extensibility. UniGame is agnostic to most UMM architectures and post-training approaches. First, since it is a general training framework that only introduces a lightweight trainable perturber, it is flexible to be integrated into most UMMs. Second, it does not conflict with existing methods, but can serve as their complement for further improvement on consistency and performance. We further explore the intergration of UniGame with emerging post-training method e.g., [jiang2025t2i, wang2024reconstruction], we train their post-trained model and demonstrate further improvements (see §4.6).
@@ 4.1 Experimental Setup
Harmon+RecA [wang2024reconstruction] 1.5B $66.94$ $0.40$ $33.67$ 56.16(+0.51)56.16~(+0.51)
@@
Baselines. We compare UniGame to two categories of baselines: (1) Different UMMs: (i) Auto-regressive models [chen2025janus, team2024chameleon, qu2025tokenflow, ge2024seed, xie2024show], which unify understanding and generation through next-token prediction in a shared token space, with improvements in vision tokenizers; (ii) Diffusion-based models [chen2025blip3, xiao2025omnigen], which leverage latent diffusion for generation while maintaining autoregressive understanding (e.g., BLIP-3o [chen2025blip3], OmniGen2 [xiao2025omnigen]); (iii) Hybrid architectures [lin2025uniworld, deng2025emerging], which employ specialized modules for different modalities (e.g., UniWorld-V1 [lin2025uniworld], BAGEL [deng2025emerging]). (2) Post-training methods: (i) Reconstruction-based alignment, which uses caption-then-reconstruct cycles to enhance understanding-generation consistency (e.g., RecA [wang2024reconstruction]); (ii) Reward-based approaches, which use reward function to refine their outputs(e.g, [jiang2025t2i]) Unlike these methods, UniGame introduces decoder-constrained self-adversarial training, converting latent inconsistencies into visually coherent counterexamples to improve reasoning robustness while preserving generation fidelity.
@@ 4.6 Extensibility and Efficiency
Extensibility. We first implement UniGame as a complement to RecA [wang2024reconstruction]. The results in Table 5 show consistent performance gains. Specifically, UniGame outperforms the original RecA by $0.5$ on MMMU for understanding and $1.27$ on UnifiedBench for consistency, while remaining the same on GenEval. These results indicate that UniGame can serve as a lightweight, plug-and-play post-training module that can be integrated into existing pipelines, requiring only minimal additional computation.
In addition to RecA, we further constructed two architectures: (1) UMM-1 uses a Qwen2.5-VL [yang2025qwen3] backbone with a SigLIP2 [zhai2023sigmoid] understanding encoder and a Stable Diffusion-1.5 [Rombach_2022_CVPR] image branch; (2) UMM-2 keeps the vision/generation stack unchanged and replaces the backbone with GPT-OSS [openai2025gptoss120bgptoss20bmodel]. Since GPT-OSS is designed for texts, we inserted a trainable 2-layer MLP that projects vision embeddings into the language space. Table 6 demonstrates that UniGame is agnostic to model architectures and can improve the performance of all three different backbones. Moreover, the gains are achieved with fewer trainable parameters (e.g., $\sim$0.45% for UMM-2 and $\sim$1.43% for UMM-1), indicating its parameter-efficient generalization.
@@
Efficiency. We further evaluate the efficiency of UniGame in comparison with RecA [wang2024reconstruction] and UAE [yan2025can] on MMMU. Table 7 shows that while achieving stronger performance, UniGame uses fewer trainable parameters, indicating its efficiency over existing post-training approaches.

### v2

@@ 1 Introduction
Unified Multimodal Models (UMMs) have recently demonstrated impressive capability in both visual understanding and image generation with a unified architecture [deng2025emerging, wang2024emu3, qu2025tokenflow, xie2024show, xiao2025omnigen, wu2025harmonizing]. By jointly leveraging a language model backbone and a visual tokenizer–decoder stack [li2025dual, chen2025janus], these models promise a unified interface for cross-modal reasoning, grounded perception, and controllable generation. Specifically, the large-scale pre-training establishes general multimodal capabilities, and the post-training stage (supervised fine-tuning, SFT, Figure 2(a)) can further improve their performance on downstream tasks with enhanced reliability.
@@
Despite their great performance, UMMs exhibit structural inconsistency between their understanding and generation pathways [yan2025can, xie2025reconstruction]. This inconsistency stems from the inherently conflicting nature of the two objectives, leading to the mismatch in various aspects such as semantics (i.e., the model can answer a question correctly yet fail to generate a corresponding image, or vice versa [Gan2020, rajabi2025token]), capability (e.g., generation is harder to improve than understanding, or vice versa [radford2021learning]), and feature compactness (e.g., understanding requires more compact feature space while generation prefers oppositely). Inconsistency widely exists in real-world applications, where models frequently encounter unexpected inputs far from the training manifold, compositional combinations unseen during training, counterfactual queries, or modality conflicts such as distribution shift [Gan2020, li2024naturalbench, oh2025understanding, oh2025vittle] and adversarial attack [li2021adversarial]. If not sufficiently studied, it would greatly undermine multimodal information fusion, model robustness, and further performance improvement.
@@
It remains challenging and unexplored to improve the consistency of UMMs, primarily due to the unclear learning objectives: only the consequences, but not the causes, are known. Therefore, recent post-training approaches tend to fill this gap using surrogate objectives. Reconstruction-based approaches (Figure 2(b)) regenerate the original images through the semantic embedding space derived from visual perception [xie2025reconstruction, yan2025can]. They optimize a unified reconstruction objective, which trains models within a closed auto-encoding loop. Reward-based methods (Figure 2(c)) typically optimize an ensembled reward function [jiang2025t2i], combining task-specific or rule-based metrics to refine the output relying on external expert models. However, both are optimized using handcrafted objectives (reconstruction or reward), which only polish model behavior on a fixed training distribution and place no explicit constraints on the two coupling branches. As a result, they reproduce behaviors within a fixed manifold rather than expanding the shared generative space, leaving the inconsistency largely unresolved.
@@ 2 Related Work
Post-training of UMMs. Aiming to resolve the inconsistency issue, existing post-training methods can be categorized into these types: (i) reconstruction and semantic-alignment losses to encourage fidelity [xie2025reconstruction, yan2025can]; (ii) RL/reward-based optimization to directly optimize downstream metrics [Ramesh2024, jiang2025t2i]. Each family improves either fidelity or robustness [Li2023, wu2024janus, xie2025reconstruction, Gan2020, rajabi2025token, Ramesh2024, qu2025tokenflow], For instance, RecA [xie2025reconstruction] leverages reconstruction alignment by conditioning generation on understanding embeddings and using reconstruction losses to bring representations closer, while T2I-R1 [jiang2025t2i] enhances image generation through collaborative semantic- and token-level chain-of-thought combined with reinforcement learning. In the vision–language domain, AT has shown potential: VILLA introduces large-scale embedding-space perturbations across image and text modalities, improving robustness and generalization [Gan2020, rajabi2025token]. Nevertheless, most methods either improve alignment or generation separately, and adversarial mechanisms are rarely incorporated into full UMMs. They commonly fail in one crucial respect: they do not exploit generation as an active adversarial process to strengthen the understanding branch. Specifically, adversarial or unguided embedding perturbations often produce off-manifold samples—unrealistic or semantically invalid artifacts, while reconstruction objectives do not intentionally surface decision-critical failure modes. Although RL-based schemes are effective, they are computationally costly and do not guarantee that these discovered examples remain decodable and semantically plausible [Ramesh2024, jiang2025t2i]. This leaves a key gap in truly improving the consistency of the understanding and generation pathways.
@@ 3.2 Motivation
Existing efforts [xie2025reconstruction, jiang2025t2i] optimize for individual goals on a fixed set of data distributions, and place no explicit constraints on the two coupling training objectives. They encourage cooperative training, reproducing existing samples rather than expanding coverage of the shared manifold, where boundary behavior is most fragile, thus aggravating inconsistency. We argue that improving consistency requires expanding this shared manifold, especially around decision boundaries, instead of merely polishing the model within its comfort regions.
@@ 3.4 The Self-Play Training Process
$\begin{split}\mathcal{L}_{U}(\theta_{U})=\;\mathbb{E}_{\mathrm{clean}}\big[\mathrm{CE}(p_{U}(\hat{a}\mid\mathbf{z},q;\theta_{U}),a)\big]+\beta\,\mathbb{E}_{\mathcal{B}}\big[\mathrm{CE}(p_{U}(\hat{a}\mid\mathbf{z},q;\theta_{U}),a)\big],\end{split}$ (6)
@@
$\begin{split}\mathcal{L}_{C}(\theta_{C};\theta_{U})=\;\mathbb{E}_{\mathrm{clean}}\mathrm{CE}\big(p_{U}(\hat{a}\mid\mathrm{Enc}(G(C(\hat{\mathbf{z}};\theta_{C}))),q;\theta_{U}),a\big)-\lambda\|\boldsymbol{\delta}\|^{2}.\end{split}$ (7)
@@ 3.5 Discussion
Extensibility. UniGame is agnostic to most UMM architectures and post-training approaches. First, since it is a general training framework that only introduces a lightweight trainable perturber, it is flexible to be integrated into most UMMs. Second, it does not conflict with existing methods, but can serve as their complement for further improvement on consistency and performance. We further explore the intergration of UniGame with emerging post-training method e.g., [jiang2025t2i, xie2025reconstruction], we train their post-trained model and demonstrate further improvements (see §4.6).
@@ 4.1 Experimental Setup
Harmon+RecA [xie2025reconstruction] 1.5B $66.94$ $0.40$ $33.67$ 56.16(+0.51)56.16~(+0.51)
@@
Baselines. We compare UniGame to two categories of baselines: (1) Different UMMs: (i) Auto-regressive models [chen2025janus, team2024chameleon, qu2025tokenflow, ge2024seed, xie2024show], which unify understanding and generation through next-token prediction in a shared token space, with improvements in vision tokenizers; (ii) Diffusion-based models [chen2025blip3, xiao2025omnigen], which leverage latent diffusion for generation while maintaining autoregressive understanding (e.g., BLIP-3o [chen2025blip3], OmniGen2 [xiao2025omnigen]); (iii) Hybrid architectures [lin2025uniworld, deng2025emerging], which employ specialized modules for different modalities (e.g., UniWorld-V1 [lin2025uniworld], BAGEL [deng2025emerging]). (2) Post-training methods: (i) Reconstruction-based alignment, which uses caption-then-reconstruct cycles to enhance understanding-generation consistency (e.g., RecA [xie2025reconstruction]); (ii) Reward-based approaches, which use reward function to refine their outputs(e.g, [jiang2025t2i]) Unlike these methods, UniGame introduces decoder-constrained self-adversarial training, converting latent inconsistencies into visually coherent counterexamples to improve reasoning robustness while preserving generation fidelity.
@@ 4.6 Extensibility and Efficiency
Extensibility. We first implement UniGame as a complement to RecA [xie2025reconstruction]. The results in Table 5 show consistent performance gains. Specifically, UniGame outperforms the original RecA by $0.5$ on MMMU for understanding and $1.27$ on UnifiedBench for consistency, while remaining the same on GenEval. These results indicate that UniGame can serve as a lightweight, plug-and-play post-training module that can be integrated into existing pipelines, requiring only minimal additional computation.
In addition to RecA, we further constructed two architectures: (1) UMM-1 uses a Qwen2.5-VL [yang2025qwen3] backbone with a SigLIP2 [zhai2023sigmoid] understanding encoder and a Stable Diffusion-1.5 [Rombach_2022_CVPR] image branch; (2) UMM-2 keeps the vision/generation stack unchanged and replaces the backbone with GPT-OSS [agarwal2025gpt]. Since GPT-OSS is designed for texts, we inserted a trainable 2-layer MLP that projects vision embeddings into the language space. Table 6 demonstrates that UniGame is agnostic to model architectures and can improve the performance of all three different backbones. Moreover, the gains are achieved with fewer trainable parameters (e.g., $\sim$0.45% for UMM-2 and $\sim$1.43% for UMM-1), indicating its parameter-efficient generalization.
@@
Efficiency. We further evaluate the efficiency of UniGame in comparison with RecA [xie2025reconstruction] and UAE [yan2025can] on MMMU. Table 7 shows that while achieving stronger performance, UniGame uses fewer trainable parameters, indicating its efficiency over existing post-training approaches.